OpenStudy (anonymous):

find the equation for y, given the derivative and the indicated point on the curve. dy/dx = 2(x-1) ; point (3,2)

OpenStudy (anonymous):

Please help??

OpenStudy (anonymous):

So is this a differential equation, or are you looking for some tangent line \(y\) to (3,2)?

OpenStudy (anonymous):

Because you can really do either.

OpenStudy (anonymous):

I am pretty sure it is a differential equation

OpenStudy (anonymous):

\[\frac{dy}{dx}=2(x-1)\] What we have here is called a "separable ordinary differential equation." By separable, we mean that, given an equation of the form \(\dfrac{dy}{dx}=f(x,y)\), we can split the function \(f\) into what I'll call component functions. These component functions are factors of \(f\), i.e. we can write \(\dfrac{dy}{dx}=g(x)h(y)\). You don't have to know all that, though. Here's what you do: \[dy=2(x-1)~dx\] Integrate both sides: \[\int dy=\int2(x-1)~dx\\ y\color{red}{+C_y}=2\left(\frac{x^2}{2}-x\color{blue}{+C_x}\right)\\ y=x^2-2x+C\] Both C's are constants, so you can combine them into one general constant \(C\). Now, given the point (3,2), you can solve for \(C\).

OpenStudy (anonymous):

Ok so now I insert 3 into the x's right?

OpenStudy (anonymous):

2 = 3^2 - 2(3) + C 2= 3 + C so C= -1?

OpenStudy (anonymous):

yes, correct